OpenStudy (anonymous):

find the dy/dx of y=3^(x^2)

OpenStudy (raden):

put ln to both sides : ln y = ln(3^(x^2))

OpenStudy (raden):

or : ln y = x^2 ln3 then take derivative : 1/y dy = 2xln3 dx dy/dx = 2xln3 * y

OpenStudy (raden):

substitute back y = 3^(x^2) so, dy/dx = 2xln3 * 3^(x^2)